# Where MLLMs Attend and What They Rely On: Explaining Autoregressive Token Generation

### v2

@@ Abstract
Multimodal large language models (MLLMs) have demonstrated remarkable capabilities in aligning visual inputs with natural language outputs. Yet, the extent to which generated tokens depend on visual modalities remains poorly understood, limiting interpretability and reliability. In this work, we present Eagle, a lightweight black-box framework for explaining autoregressive token generation in MLLMs. Eagle attributes any selected tokens to compact perceptual regions while quantifying the relative influence of language priors and perceptual evidence. The framework introduces an objective function that unifies sufficiency (insight score) and indispensability (necessity score), optimized via greedy search over sparsified image regions for faithful and efficient attribution. Beyond spatial attribution, Eagle performs modality-aware analysis that disentangles what tokens rely on, providing fine-grained interpretability of model decisions. Extensive experiments across open-source MLLMs show that Eagle consistently outperforms existing methods in faithfulness, localization, and hallucination diagnosis, while requiring substantially less GPU memory. These results highlight its effectiveness and practicality for advancing the interpretability of MLLMs. The code will be released at https://ruoyuchen10.github.io/EAGLE/.
@@ 1 Introduction
Multimodal large language models (MLLMs) (Achiam et al., 2023; Wang et al., 2025; Bai et al., 2025; Comanici et al., 2025) have achieved significant progress in vision–language understanding and generation. By jointly modeling visual and textual modalities, they can now perform a wide range of tasks, such as image captioning and visual question answering (VQA) (Li et al., 2025b). These advances have enabled MLLMs to approach human-level performance on many benchmarks and to underpin various real-world applications (Liang et al., 2024; Li et al., 2024). However, alongside these advances come critical challenges in transparency and reliability (Zhang et al., 2025b). As parameter scales and modality coverage continue to expand, MLLMs become increasingly opaque, making it difficult to trace how specific inputs influence generated outputs (Xing et al., 2025; Chen et al., 2025c; b). Furthermore, MLLMs are susceptible to hallucinations (Chen et al., 2025b; a), which undermine trust in safety-critical domains such as healthcare (Ahmed et al., 2025) and autonomous driving (Chen et al., 2024a). These limitations highlight the urgent need for efficient and faithful attribution methods to improve decision transparency, diagnose errors, and enhance the safety and trustworthiness of MLLMs (Lin et al., 2025; Dang et al., 2024; Liang et al., 2025b; 2023; a; Lu et al., 2025).
@@
Attribution in MLLMs is particularly challenging because they generate tokens autoregressively, making classification-based attribution methods difficult to adapt. Attention visualization approaches (Ben Melech Stan et al., 2024) often fail to capture complex cross-modal interactions, while gradient-based extensions (Zhang et al., 2025b; Xing et al., 2025) aggregate token logits but remain confounded by textual priors. More recently, TAM (Li et al., 2025a) employed activation maps to explain individual tokens and showed promising localization on Qwen2-VL (Wang et al., 2024), yet it cannot generalize to all MLLMs or capture multi-token contributions. In summary, attribution methods based on activation maps or gradients face inherent limitations: (1) activation-based approaches lack a direct causal link between inputs and outputs, reflecting only intermediate layer preferences often misaligned with human intuition; and (2) gradient-based approaches are sensitive to cumulative effects in long sequences and easily disturbed by noise and modality imbalance.
@@
To more faithfully explain the generation of MLLMs, we propose Eagle (Explaining Autoregressive Generation by Language priors or Evidence), a black-box attribution framework for interpreting autoregressive token generation. As shown in Fig 1, our method supports attribution for any chosen set of output tokens, revealing the perceptual regions that drive their generation and quantifying the relative roles of language priors and visual evidence. Inspired by submodular subset selection, we aim to find the minimal set of perceptual regions that maximizes token logits, conditioned on the prompt and context. We design an objective function with two components: the insight score, capturing regions sufficient for generation, and the necessity score, identifying regions whose removal impairs generation. By applying greedy search over sparsified image regions, we construct an ordered ranking that attributes which perceptual regions promote generation in MLLMs, addressing the question of “Where MLLMs Attend”. Beyond spatial attribution, we also assess “What They Rely On”. By tracking how token logits evolve as salient regions are progressively introduced, we measure whether each token depends more on perceptual evidence or language priors, offering a faithful and comprehensive view of model decisions.
@@
We evaluate our method on open-source MLLMs, including LLaVA-1.5 (Liu et al., 2024), Qwen2.5-VL (Bai et al., 2025), and InternVL3.5 (Wang et al., 2025), using the MS COCO Lin et al. (2014) and MMVP (Tong et al., 2024) datasets for image captioning and VQA. On faithfulness metrics, our approach outperforms existing attribution methods (LLaVA-CAM (Zhang et al., 2025b), IGOS++ (Xing et al., 2025), and TAM (Li et al., 2025a)) by an average of 20.0% in insertion and 13.4% in deletion for image captioning, and by 20.6% and 8.1% on the same metrics for VQA. At the word level, our method achieves more rational explanations of object tokens, surpassing TAM by 36.42% and 42.63% on the Pointing Game under box-level and mask-level annotations, respectively. Finally, on the RePOPE benchmark Neuhaus & Hein (2025) for object hallucination, our method accurately localizes the visual elements responsible for hallucinations and mitigates them by removing only a minimal set of interfering regions. These results demonstrate the versatility of our method across diverse tasks and benchmarks.
@@
    We propose Eagle, a lightweight black-box attribution framework for autoregressive token generation, which attributes any selected set of tokens to compact perceptual regions with low GPU memory cost.
@@ 2 Related Work
Multimodal LLMs Attribution. Research on input-level attribution for Multimodal Large Language Models (MLLMs) is still nascent. LVLM-Interpret (Ben Melech Stan et al., 2024) visualizes alignment between LLaVA outputs and images using raw attention, while LLaVA-CAM (Zhang et al., 2025b) adapts Smooth-CAM (Omeiza et al., 2019) to token-level probabilities, but both suffer from layer sensitivity and limited faithfulness. VPS (Chen et al., 2025b) introduces a search-based method for object-level tasks, yet it is restricted to grounding and detection. IGOS++(Xing et al., 2025) identifies visually aligned tokens but remains parameter-sensitive. More recently, TAM(Li et al., 2025a) reduces contextual noise in activation maps, improving token-level attribution. However, gradient-based methods remain memory-intensive and unstable. In contrast, we propose a black-box attribution framework that localizes outputs to compact input regions without relying on token selection, quantifies the influence of language priors versus perceptual evidence, and further explains the causes of object hallucinations in MLLMs.
Interpreting Hallucinations in MLLMs. Several studies have applied interpretability techniques to examine hallucinations. Jiang et al. (2025) investigated how image latent representations in vision-language models are projected into the language vocabulary, thereby shaping the model’s confidence in both “real” and “hallucinatory” objects, and further proposed a representation correction method to mitigate hallucinations. Zhang et al. (2025a) examined whether MLLMs attend to incorrect regions when producing wrong answers, leveraging their internal attention maps. VaLSe (Chen et al., 2025a) employs gradient- and attention-based attribution maps to identify noisy regions that contribute to hallucinations. In this work, we primarily focus on interpreting which input regions lead to incorrect decisions, aiming to suppress hallucinations by removing as few regions as possible.
@@ 3.1 Task Formulation
For interpretability analysis, our objective is to identify the image regions $\mathbf{x}$ that most strongly drive the model’s decisions. Image features in MLLMs are typically high-dimensional and information-dense but also redundant and less directly interpretable than text. We therefore focus on decomposing $\mathbf{x}$ into semantically meaningful subregions. Specifically, the image is sparsified into $V=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}$ using the SLICO (Achanta et al., 2012) superpixel segmentation method, where $\mathbf{x}_{i}$ denotes the $i$-th subregion. The attribution problem is then cast as a subset selection task (Chen et al., 2024b): $\max_{S\subseteq V,|S|<k}\mathcal{F}(S)$, where $k$ is the maximum number of selected subregions and $\mathcal{F}(\cdot)$ is a set function measuring interpretability. Beyond the unordered case, attribution also depends on the order in which regions contribute to the decision. We therefore extend the formulation to ordered subsets:
@@ 3.2 Explaining Autoregressive Generation
$s_{\text{insight}}(S,\texttt{Prompt},\mathbf{y},T,\mathcal{V})=\sum_{i=1}^{n}p\!\left(y_{t_{i}}=v_{i}\mid S,\texttt{Prompt},\mathbf{y}_{<t_{i}}\right),$ (2)
@@
Necessity Score: Another key metric for interpretability is the identification of the minimal set of input regions whose removal leads to a significant decrease in the probability of generating the target label, thereby revealing the indispensable evidence that the model relies on. Formally, for a candidate subregion $S$, the necessity score is defined as:
$s_{\text{necessity}}(V\setminus S,\texttt{Prompt},\mathbf{y},T,\mathcal{V})=\sum_{i=1}^{n}\Big(1-p\!\left(y_{t_{i}}=v_{i}\mid V\setminus S,\texttt{Prompt},\mathbf{y}_{<t_{i}}\right)\Big),$ (3)
@@
$\mathcal{F}(S,V,\texttt{Prompt},\mathbf{y},T,\mathcal{V})=s_{\text{insight}}(S,\texttt{Prompt},\mathbf{y},T,\mathcal{V})+s_{\text{necessity}}(V\setminus S,\texttt{Prompt},\mathbf{y},T,\mathcal{V}),$ (4)
@@
Saliency Map Generation: To optimize the objective in Eq. 1, an $\mathcal{NP}$-hard problem, we adopt a greedy search strategy. At each step, the region yielding the largest marginal gain with respect to the objective function is added to the current set until the budget $k$ is reached, producing an ordered set $\pi$. Beyond ranking, it is also important to assess the relative saliency differences among subregions. We evaluate these differences by examining the marginal gains of the objective function as the ordered subset expands. A larger gain indicates that the newly added subregion remains highly influential, whereas diminishing gains approaching zero suggest that subsequent subregions contribute negligibly and exhibit limited saliency distinction. The attribution score $\mathcal{A}_{i}$ for a subregion $\pi_{i}$ within the ordered set $\pi$ is defined as:
$\mathcal{A}_{i}=\begin{cases}0&\text{if }i=1,\\ \mathcal{A}_{i-1}-\big|\mathcal{F}(\pi_{:i})-\mathcal{F}(\pi_{:i-1})\big|&\text{if }i>1,\end{cases}$ (5)
where $\pi_{:i}$ denotes the combination of the top $i$ subregions, and the attribution scores start from zero, decrease progressively with each step, and are subsequently normalized.
@@
Beyond identifying which perceptual regions promote the generation of specific autoregressive tokens, we further analyze whether each generated token is more strongly influenced by language priors or by perceptual evidence. Existing approaches often assess token relevance to the visual modality by observing changes in probability when the input image is masked Xing et al. (2025). However, simply comparing the probability with the full image against that without the image is not a reliable indicator of visual relevance, as the probability may first increase and then decrease when visual inputs are progressively inserted (Chen et al., 2024b). By contrast, if a token is truly irrelevant to the visual modality, its probability should remain stable regardless of how the image is modified.
@@
$I_{t_{i}}=\sum_{r=1}^{|\pi|}\Big(p\!\left(y_{t_{i}}=v_{i}\mid\pi_{:r},\texttt{Prompt},\mathbf{y}_{<t_{i}}\right)-\min_{1\leq j\leq|\pi|}p\!\left(y_{t_{i}}=v_{i}\mid\pi_{:j},\texttt{Prompt},\mathbf{y}_{<t_{i}}\right)\Big),$ (6)
where $v_{i}$ denotes the vocabulary index of the target token $y_{t_{i}}$. The influence score $I_{t_{i}}$ measures the impact of perceptual evidence on the generation of token $y_{t_{i}}$. A larger score indicates that the token generation is more strongly driven by perceptual evidence, whereas a smaller score suggests a greater reliance on language priors, as shown in Fig. 2. The detailed calculation process of the proposed Eagle algorithm is outlined in Algorithm 1.
Remark 1 (Weak Submodularity).
Our objective function $\mathcal{F}(\cdot)$ is not strictly submodular in MLLMs. However, it exhibits weak submodularity, a relaxed condition that bounds the deviation from true submodularity. Formally, let $\gamma\in(0,1]$ denote the submodularity ratio of $\mathcal{F}$:
$\gamma=\min_{L\subseteq U,S\subseteq U\setminus L}\frac{\sum_{i\in S}\big(\mathcal{F}(L\cup\{i\})-\mathcal{F}(L)\big)}{\mathcal{F}(L\cup S)-\mathcal{F}(L)}.$
When $\gamma=1$, $\mathcal{F}$ is strictly submodular; smaller $\gamma$ values indicate weaker submodularity. Under weak submodularity, greedy selection is still guaranteed to achieve a $(1-e^{-\gamma})$-approximation of the optimal solution (Bian et al., 2017). Thus, the stronger the submodular property of $\mathcal{F}$ in MLLMs (i.e., larger $\gamma$), the tighter the theoretical bound and the more reliable the approximation.
@@
Gradient-based methods (Xing et al., 2025) rely on selecting visually relevant tokens; choosing tokens dominated by language priors can distort attribution and yield unreliable explanations. In contrast, our approach is token-agnostic: even when applied to tokens strongly influenced by language priors, the visual attribution remains unaffected. Moreover, after attribution, our framework explicitly evaluates whether the selected tokens are primarily driven by perceptual evidence or language priors.
Remark 3 (Interactive Token-Level Explanation).

### v4

@@ Abstract
Multimodal large language models (MLLMs) have demonstrated remarkable capabilities in aligning visual inputs with natural language outputs. Yet, the extent to which generated tokens depend on visual modalities remains poorly understood, limiting interpretability and reliability. In this work, we present Eagle, a lightweight black-box framework for explaining autoregressive token generation in MLLMs. Eagle attributes any selected tokens to compact perceptual regions while quantifying the relative influence of language priors and perceptual evidence. The framework introduces an objective function that unifies sufficiency (insight score) and indispensability (necessity score), optimized via greedy search over sparsified image regions for faithful and efficient attribution. Beyond spatial attribution, Eagle performs modality-aware analysis that disentangles what tokens rely on, providing fine-grained interpretability of model decisions. Extensive experiments across open-source MLLMs show that Eagle consistently outperforms existing methods in faithfulness, localization, and hallucination diagnosis, while requiring substantially less GPU memory. These results highlight its effectiveness and practicality for advancing the interpretability of MLLMs.
@@ 1 Introduction
Multimodal large language models (MLLMs) [2, 32, 4, 12] have achieved significant progress in vision–language understanding and generation. By jointly modeling visual and textual modalities, they can now perform a wide range of tasks, such as image captioning and visual question answering (VQA) [18]. These advances have enabled MLLMs to approach human-level performance on many benchmarks and to underpin various real-world applications [19, 16]. However, alongside these advances come critical challenges in transparency and reliability [36]. As parameter scales and modality coverage continue to expand, MLLMs become increasingly opaque, making it difficult to trace how specific inputs influence generated outputs [33, 9, 8]. Furthermore, MLLMs are susceptible to hallucinations [8, 6], which undermine trust in safety-critical domains such as healthcare [3] and autonomous driving [7]. These limitations highlight the urgent need for efficient and faithful attribution methods to improve decision transparency, diagnose errors, and enhance the safety and trustworthiness of MLLMs [24, 13, 21, 22, 20, 26].
@@
Attribution in MLLMs is particularly challenging because they generate tokens autoregressively, making classification-based attribution methods difficult to adapt. Attention visualization approaches [5] often fail to capture complex cross-modal interactions, while gradient-based extensions [36, 33] aggregate token logits but remain confounded by textual priors. More recently, TAM [17] employed activation maps to explain individual tokens and showed promising localization on Qwen2-VL [31], yet it cannot generalize to all MLLMs or capture multi-token contributions. In summary, attribution methods based on activation maps or gradients face inherent limitations: (1) activation-based approaches lack a direct causal link between inputs and outputs, reflecting only intermediate layer preferences often misaligned with human intuition; and (2) gradient-based approaches are sensitive to cumulative effects in long sequences and easily disturbed by noise and modality imbalance. The subset-selection–based VPS algorithm [8] has achieved notable advances in interpreting visual grounding, outperforming gradient- and activation-based methods. However, its objective function cannot be directly transferred to MLLMs.
@@
To more faithfully explain the generation of MLLMs, we propose Eagle (Explaining Autoregressive Generation by Language priors or Evidence), a black-box attribution framework for interpreting autoregressive token generation. As shown in Fig 1, our method supports attribution for any chosen set of output tokens, revealing the perceptual regions that drive their generation and quantifying the relative roles of language priors and visual evidence. Inspired by VPS [8], a submodular subset selection based attribution method, we aim to find the minimal set of perceptual regions that maximizes token logits, conditioned on the prompt and context. We design an objective function with two components special for MLLMs: the insight score, capturing regions sufficient for generation, and the necessity score, identifying regions whose removal impairs generation. By applying greedy search over sparsified image regions, we construct an ordered ranking that attributes which perceptual regions promote generation in MLLMs, addressing the question of “Where MLLMs Attend”. Beyond spatial attribution, we also assess “What They Rely On”. By tracking how token logits evolve as salient regions are progressively introduced, we measure whether each token depends more on perceptual evidence or language priors, offering a faithful and comprehensive view of model decisions.
@@
We evaluate our method on open-source MLLMs, including LLaVA-1.5 [25], Qwen2.5-VL [4], and InternVL3.5 [32], using the MS COCO [23] and MMVP [30] datasets for image captioning and VQA. On faithfulness metrics, our approach outperforms existing attribution methods (LLaVA-CAM [36], IGOS++ [33], and TAM [17]) by an average of 20.0% in insertion and 13.4% in deletion for image captioning, and by 20.6% and 8.1% on the same metrics for VQA. At the word level, our method achieves more rational explanations of object tokens, surpassing TAM by 36.42% and 42.63% on the Pointing Game under box-level and mask-level annotations, respectively. Finally, on the RePOPE benchmark [27] for object hallucination, our method accurately localizes the visual elements responsible for hallucinations and mitigates them by removing only a minimal set of interfering regions. These results demonstrate the versatility of our method across diverse tasks and benchmarks.
@@
    We propose Eagle, a lightweight black-box attribution framework for autoregressive token generation, which attributes any selected set of tokens to compact perceptual regions with low GPU memory cost. It further reveals the latent potential of attribution performance in MLLMs.
@@ 2 Related Work
Multimodal LLMs Attribution. Research on input-level attribution for Multimodal Large Language Models (MLLMs) is still nascent. LVLM-Interpret [5] visualizes alignment between LLaVA outputs and images using raw attention, while LLaVA-CAM [36] adapts Smooth-CAM [28] to token-level probabilities, but both suffer from layer sensitivity and limited faithfulness. VPS [8] introduces a search-based method for object-level tasks, yet it is restricted to grounding and detection. IGOS++[33] identifies visually aligned tokens but remains parameter-sensitive. More recently, TAM[17] reduces contextual noise in activation maps, improving token-level attribution. However, gradient-based methods remain memory-intensive and unstable. In contrast, we propose a black-box attribution framework that localizes outputs to compact input regions without relying on token selection, quantifies the influence of language priors versus perceptual evidence, and further explains the causes of object hallucinations in MLLMs.
Interpreting Hallucinations in MLLMs. Several studies have applied interpretability techniques to examine hallucinations. [15] investigated how image latent representations in vision-language models are projected into the language vocabulary, thereby shaping the model’s confidence in both “real” and “hallucinatory” objects, and further proposed a representation correction method to mitigate hallucinations. [35] examined whether MLLMs attend to incorrect regions when producing wrong answers, leveraging their internal attention maps. VaLSe [6] employs gradient- and attention-based attribution maps to identify noisy regions that contribute to hallucinations. In this work, we primarily focus on interpreting which input regions lead to incorrect decisions, aiming to suppress hallucinations by removing as few regions as possible.
@@ 3.1 Task Formulation
For interpretability analysis, inspired by VPS [8], our objective is to identify the image regions $\mathbf{x}$ that most strongly drive the model’s decisions. Image features in MLLMs are typically high-dimensional and information-dense but also redundant and less directly interpretable than text. We therefore focus on decomposing $\mathbf{x}$ into semantically meaningful subregions. Specifically, the image is sparsified into $V=\{\mathbf{x}_{1},\mathbf{x}_{2},\dots,\mathbf{x}_{N}\}$ using the SLICO [1] superpixel segmentation method, where $\mathbf{x}_{i}$ denotes the $i$-th subregion. The attribution problem is then cast as a subset selection task [10]: $\max_{S\subseteq V,|S|<k}\mathcal{F}(S)$, where $k$ is the maximum number of selected subregions and $\mathcal{F}(\cdot)$ is a set function measuring interpretability. Beyond the unordered case, attribution also depends on the order in which regions contribute to the decision. We therefore extend the formulation to ordered subsets:
@@ 3.2 Explaining Autoregressive Generation
$s_{\text{insight}}$ $(S,\texttt{Prompt},\mathbf{y},T,\mathcal{V})=$ (2)
$\sum_{i=1}^{n}p\!\left(y_{t_{i}}=v_{i}\mid S,\texttt{Prompt},\mathbf{y}_{<t_{i}}\right),$
@@
Necessity Score: Another key metric for interpretability is the identification of the minimal set of input regions whose removal leads to a significant decrease in the probability of generating the target label, thereby revealing the indispensable evidence that the model relies on. Formally, the necessity score is defined as:
$s_{\text{necessity}}$ $(V\setminus S,\texttt{Prompt},\mathbf{y},T,\mathcal{V})=$ (3)
$\sum_{i=1}^{n}\Big(1-p\!\left(y_{t_{i}}=v_{i}\mid V\setminus S,\texttt{Prompt},\mathbf{y}_{<t_{i}}\right)\Big),$
@@
$\mathcal{F}(S,V,\texttt{Prompt},$ $\mathbf{y},T,\mathcal{V})=s_{\text{insight}}(S,\texttt{Prompt},\mathbf{y},T,\mathcal{V})$ (4)
$+s_{\text{necessity}}(V\setminus S,\texttt{Prompt},\mathbf{y},T,\mathcal{V}),$
@@
Saliency Map Generation: Similar to the calculation process for VPS [8], to optimize the objective in Eq. 1, an $\mathcal{NP}$-hard problem, a greedy search strategy is adopted. Following [8], we assess saliency contrast across subregions via marginal increments of the objective evaluated along the ordered sequence. Subregions yielding larger increments are deemed influential; diminishing increments approaching zero suggest that later subregions add little information and exhibit weak saliency differentiation.
@@
Beyond identifying which perceptual regions promote the generation of specific autoregressive tokens, we further analyze whether each generated token is more strongly influenced by language priors or by perceptual evidence. Existing approaches often assess token relevance to the visual modality by observing changes in probability when the input image is masked [33]. However, simply comparing the probability with the full image against that without the image is not a reliable indicator of visual relevance, as the probability may first increase and then decrease when visual inputs are progressively inserted [10]. By contrast, if a token is truly irrelevant to the visual modality, its probability should remain stable regardless of how the image is modified.
@@
$I_{t_{i}}$ $=\sum_{r=1}^{|\pi|}\Big(p\!\left(y_{t_{i}}=v_{i}\mid\pi_{:r},\texttt{Prompt},\mathbf{y}_{<t_{i}}\right)$ (5)
$\quad-\min_{1\leq j\leq|\pi|}p\!\left(y_{t_{i}}=v_{i}\mid\pi_{:j},\texttt{Prompt},\mathbf{y}_{<t_{i}}\right)\Big).$
where $v_{i}$ denotes the vocabulary index of the target token $y_{t_{i}}$. The influence score $I_{t_{i}}$ measures the impact of perceptual evidence on the generation of token $y_{t_{i}}$. A larger score indicates that the token generation is more strongly driven by perceptual evidence, whereas a smaller score suggests a greater reliance on language priors, as shown in Fig. 2. The detailed calculation process of the proposed Eagle algorithm is outlined in the supplementary materials
@@
Gradient-based methods [33] rely on selecting visually relevant tokens; choosing tokens dominated by language priors can distort attribution and yield unreliable explanations. In contrast, our approach is token-agnostic: even when applied to tokens strongly influenced by language priors, the visual attribution remains unaffected. Moreover, after attribution, our framework explicitly evaluates whether the selected tokens are primarily driven by perceptual evidence or language priors.
Remark 2 (Interactive Explanation).
@@
Remark 4 (Optimality of the Greedy Search).
The problem reduces to maximizing a monotone submodular function under a cardinality constraint, where greedy yields the tight (11/e)(1-1/e)-approximation and is essentially optimal unless $P=NP$ [14]. Thus, to probe the upper bound of attribution quality, greedy search is used without applying acceleration techniques.